### v1

@@ 3.3 Co-moving number density of SMBHs
Next, assuming the detected variable sources are AGN powered by accreting SMBHs, we use the numbers of variable sources to estimate the co-moving number density of SMBHs, $n_{\rm SMBH}(z)$. As discussed in Paper I, we expect variable sources detected at high redshifts to be more likely to be AGN rather than supernovae or other stellar transients. Also, our focus on galactic nuclei will also tend to reduce supernova interlopers in our AGN sample.
@@
We consider $n_{\rm SMBH,raw-fp}$ to be the most reliable, direct estimate of a lower limit to the co-moving number density of SMBHs. These values are plotted as open squares in Figure 5. We see that in the lowest redshift bin from $z=0$ to 0.5 that $n_{\rm SMBH,raw-fp}\simeq 3\times 10^{-2}\>{\rm cMpc}^{-3}$ for the $2\sigma$ selected sample, dropping to $\simeq 10^{-2}\>{\rm cMpc}^{-3}$ for the $3\sigma$ sources. We note that these values are a few times higher than the previous $z\sim 0$ estimate of $n_{\rm SMBH}$ estimated by Vika et al. (2009) and Banik et al. (2019). At the highest redshifts, i.e., $z\sim 6-9$, we find $n_{\rm SMBH,raw-fp}\gtrsim 10^{-4}\>{\rm cMpc}^{-3}$. In between these redshift extremes we derive smoothly declining intermediate values of $n_{\rm SMBH,raw-fp}$ as redshift increases.
@@
After the above correction for variability incompleteness, in the low-$z$ interval we obtain co-moving number densities of SMBHs of $n_{\rm SMBH,var}\simeq 6\times 10^{-2}\>{\rm cMpc}^{-3}$ (open diamonds in Fig. 5), with little sensitivity to the choice of significance level. This may indicate our estimate of the variability incompleteness factor as a function of significance level is reasonable. The smallest Poisson uncertainties result from the $2\sigma$ selected sample, formally yielding $n_{\rm SMBH,var}=6.5\pm 2.9\times 10^{-2}\>{\rm cMpc}^{-3}$. We regard this estimate as our best measure of the true local number density of SMBHs. However, it is likely to still be a lower limit if there are very faint AGN whose flux and flux variations would not be detected by the HST observations. At the highest redshifts, $z=7-9$, we find $n_{\rm SMBH,var}\simeq 5\times 10^{-4}\>{\rm cMpc}^{-3}$. Given that we have adopted a redshift independent variability correction factor, the redshift dependence of $n_{\rm SMBH,var}$ is the same as that of $n_{\rm SMBH,raw-fp}$.
@@
Finally, we correct for “luminosity incompleteness” via a luminosity correction factor, $F_{\rm lum}$, i.e., $n_{\rm SMBH,lum}=F_{\rm lum}n_{\rm SMBH,var}$. In particular, at high redshift we are only able to detect relatively bright AGN (e.g., at $z\sim 6-7$ this corresponds at $M_{\rm UV}\simeq-18.6$), so this correction factor will boost the inferred number density of SMBHs. Following the method of Paper I and Harikane et al. (2023), who detected AGN via broad emission lines, we extrapolate from our detected AGN down an assumed luminosity function to a level of $M_{\rm UV}=-17\>$mag. We note that this is an arbitrary level, but by adopting this choice we are able to make a fair comparison to these previous studies. The value of the luminosity correction factor, $F_{\rm lum}$, generally becomes larger with redshift, but also depends on the form of the UV luminosity function (UVLF) we assume. Specifically, we integrate the double power-law fits detailed in Finkelstein & Bagley (2022) from redshift $z\sim 3$ to 9. At lower redshift, we integrate Schechter function fits to the UVLF from UVCANDELS and GALEX survey programs (Arnouts et al., 2005; Sun et al., 2024). At $z\lesssim 4$ we see that $F_{\rm lum}$ becomes smaller than unity as our detection limit becomes fainter than the UVLF lower bound of $M_{\rm UV}=-17$. Note that in the low $z$ interval, i.e., at $z=0-0.5$, our sensitivity corresponds to an absolute magnitude of about $M_{\rm UV}=-8.5$ assuming an average redshift of $z\sim 0.25$. This explains why we consider $n_{\rm SMBH,var}$ as our best estimate at low redshifts. In Table 4 we report the values of $F_{\rm lum}$ used in each redshift interval.
@@
After the above correction factor for luminosity incompleteness, in the high-redshift regime ($z=6-9$) we find $n_{\rm SMBH,lum}\sim 10^{-2}\>{\rm cMpc}^{-3}$. The value with the smallest error bar is derived from the $2\sigma$ sample in the interval $z=7-9$, i.e., $n_{\rm SMBH,lum}\simeq 1.5\pm 1.0\times 10^{-2}\>{\rm cMpc}^{-3}$. We caution that this result involves a relatively large luminosity function correction factor of $F_{\rm lum}\simeq 18$ and that the systematic uncertainty associated with this correction factor is not included in the above estimate. The redshift 6 to 7 interval has similarly high values of $n_{\rm SMBH,lum}$ (although the sample based on $2\sigma$ significance does not yield a number of sources greater than those expected from false positives). At intermediate redshifts, $z=3-6$ we notice that $n_{\rm SMBH,lum}$ takes smaller values $\sim 10^{-3}\>{\rm cMpc}^{-3}$. We do not expect such a rapid decline in the true number density of SMBHs, so this result may either indicate that a larger fraction of SMBHs were brighter than $M_{\rm UV}=-17\>$mag at $z>6$ or that the values at high redshift are somewhat overestimated via their luminosity function correction factor. At low redshifts, we see $n_{\rm SMBH,lum}$ rises again to values of a few ×$\times 10^{-3}\>{\rm cMpc}^{-3}$. However, at $z=0-0.5$ this is a factor of about 20 smaller than the number we infer after variability completeness correction. This is simply a reflection of the fact that the observations are able to detect AGN that are much fainter than the arbitrary limit of $M_{\rm UV}=-17\>$mag.
@@
In Figure 6 we show our $3\sigma$ results for $n_{\rm SMBH,var}$ and $n_{\rm SMBH,lum}$, along with several other estimates of $n_{\rm SMBH}$. As mentioned, the previous estimate of $n_{\rm SMBH}(z=0)$ of $\sim 5\times 10^{-3}\>{\rm cMpc}^{-3}$ (Banik et al., 2019) and $\sim 9\times 10^{-3}\>{\rm cMpc}^{-3}$ (Vika et al., 2009) now appear to be superseded by about a factor of 10 with our HUDF variability study estimate of $n_{\rm SMBH,var}$. Similarly, we see that our values of $n_{\rm SMBH,var}$ out to $z=3$ are about 10 times higher than the values of $n_{\rm SMBH}$ implied by the 31 previously known AGN in the HUDF (Lyu et al., 2022).
@@
At high redshifts we find consistency with our result from Paper I based on three detected AGN candidates between $z=6$ and 7. Our results for $n_{\rm SMBH,lum}$ at $z>6$ exceed those inferred from the JWST selected sample of Harikane et al. (2023) by about a factor of 10 and extend to even higher redshifts up to $\sim 9$. From $z=4$ to 6 our estimates of $n_{\rm SMBH,lum}$ are similar to those of Harikane et al. (2023). For all these high-$z$ results, there are large uncertainties due to the small number of directly detected sources. Indeed our $2\sigma$ selected sample at $z>7$ has the smallest Poisson errors.
@@
$>2\sigma$ 247 219 218 596 15/31 2.07
$>2.5\sigma$ 112 74 83 233 11/31 2.81
$>3\sigma$ 54 41 39 106 10/31 3.10
$>3\sigma$ + diff. imaging 60 50 42 124 10/31 3.10
@@
Figure 6: As Fig. 5, but now only showing $n_{\rm SMBH}$ estimated from $3\sigma$ variables. Various other observational constraints are displayed: the red diamond shows the observational constraint derived in Paper I (Hayes et al., 2024) based on three detected AGN between $z=6$ and 7. Broad emission line sources at $z=4-7$ (Harikane et al., 2023) are shown with the black diamonds, which use similar luminosity correction factors as our study. Brown stars indicate the number densities from the sample of 31 previously known AGN in the HUDF (Lyu et al., 2022). The black square shows a previous estimate of the local ($z=0$) co-moving number density of SMBHs (Banik et al., 2019).
@@ 3.4 Implications for SMBH Seeding Models
Our estimates of $n_{\rm SMBH}$ shown in Figures 5 and 6 place strong constraints on SMBH seeding models. At high redshifts, $z>6$, we see that our estimates of $n_{\rm SMBH,lum}\gtrsim 10^{-2}\>{\rm cMpc}^{-2}$ are about a factor of 50 higher than the fiducial Halo Mass Threshold (HMT) model used in the Illustris family of simulations (Vogelsberger et al., 2014) in which dark matter halos with masses $>7\times 10^{10}\>M_{\odot}$ are seeded with SMBHs. The range of the green band shown in Figs. 5 and 6 corresponds to varying this threshold mass by factors of two. Thus an HMT model with a threshold mass of $\sim 3\times 10^{10}\>M_{\odot}$, which would be consistent with the $z=0$ constraints, would be closer to the high-$z$ data, but still low by about a factor of 10.
Figures 5 and 6 also compares our constraints with the Romulus25 simulation (Tremmel et al., 2017). We see that this simulation, which seeds SMBHs using simple threshold conditions of gas properties (see §1), produces a high number density of SMBHs at early times, i.e., $n_{\rm SMBH}\sim 5\times 10^{-2}\>{\rm cMpc}^{-3}$ at $z\sim 12$. This rises by almost an order of magnitude by $z\simeq 4$, before then undergoing a modest decline by $z=0$. Given its high number densities of SMBHs, this model remains consistent with our observational constraints on $n_{\rm SMBH}$.
As discussed in §1, Direct Collapse (DC) seeing models struggle to reproduce the overall number densities of SMBHs. The most optimistic of the DC models discussed in §1 is that of Chon et al. (2016) with $n_{\rm SMBH}\sim 10^{-4}\>{\rm Mpc}^{-3}$. As shown in Figs. 5 and 6 this is about a factor of 100 below $n_{\rm SMBH}$ inferred from our HUDF variability study. Furthermore, the later simulations of DC by Wise et al. (2019) and Latif et al. (2022) find values of $n_{\rm SMBH}\lesssim 10^{-6}\>{\rm Mpc}^{-3}$, i.e., at least a factor of $10^{4}$ below our high-$z$ estimate. The constraints are even more severe when considering our low-$z$ estimates of $n_{\rm SMBH}$, which have smaller uncertainties and values of $\simeq 5\times 10^{-2}\>{\rm Mpc}^{-3}$.
@@
where $t_{*,10}=t_{*}/10\>$Myr, $t_{*}$ is the lifetime of the supermassive star, $S_{53}\equiv S/10^{53}\>{\rm s}^{-1}$, $S$ is the H-ionizing luminosity of the star, and $R_{R}$ is the radius of R-type HII region expansion achieved in the intergalactic medium. This estimate can be viewed as an upper limit since it has assumed maximal close packing of the sources and ignored other contributions to the ionizing background from Pop III.2, Pop II and AGN sources. It is interesting that our estimate of $n_{\rm SMBH}$ at low-$z$ from the HUDF is within a factor of a few of this fiducial estimate.
@@ 3.5 Comparison with other AGN diagnostics
Despite recent advances, it remains immensely challenging to obtain an accurate census of AGN at $z\gtrsim 6$ at faint magnitudes. Hot dust tracers (e.g., Stern et al., 2005) are shifted to the FIR, and X-ray and radio facilities are not sufficiently sensitive to identify individual faint AGN at these distances. High ionization UV emission lines (e.g., He ii $\lambda$1640, iv $\lambda$1550, v $\lambda$1240) are sometimes observed in luminous targets (e.g., Mainali et al., 2017; Laporte et al., 2017), but they can be inconclusive, as these lines also form in low-metallicity starbursts (Senchyna et al., 2020; Berg et al., 2019; Saxena et al., 2020). JWST-NIRSpec micro-shutter array (MSA) observations have delivered deep IR spectroscopy and several studies (e.g., Cameron et al., 2023; Boyett et al., 2024; Hu et al., 2024) could identify AGN via line flux ratios classification (BPT-like diagrams, Baldwin et al., 1981), although the targeted approach of MSA spectroscopy is subject to photometric pre-selection biases and tends to be focused more towards the brighter end of the luminosity function. In addition, NIRSpec surveys face challenges in achieving completeness at $M_{UV}\simeq-17$ to $-20$, where galaxies are $\sim 100$ times more abundant.

### v2

@@ 3.3 Co-moving number density of SMBHs
Next, assuming the detected variable sources are AGN powered by accreting SMBHs, we use the numbers of variable sources to estimate the co-moving number density of SMBHs, $n_{\rm SMBH}(z)$. As discussed in Paper I, we expect variable sources detected at high redshifts to be more likely to be AGN rather than supernovae or other stellar transients. Also, our focus on galactic nuclei will also tend to reduce supernova interlopers in our AGN sample. Very high redshift sources, i.e. $z\gtrsim 6$, where the lower 1 sigma uncertainty in redshift is compatible with a low redshift interloper scenario, i.e., $z_{phot}-z_{low}\gtrsim 2$, are discarded from the number density estimate.
@@
We consider $n_{\rm SMBH,raw-fp}$ to be the most reliable, direct estimate of a lower limit to the co-moving number density of SMBHs. These values are plotted as open squares in Figure 6. We see that in the lowest redshift bin from $z=0$ to 0.5 that $n_{\rm SMBH,raw-fp}\simeq 6\times 10^{-3}\>{\rm cMpc}^{-3}$ regardless of the $\sigma$ threshold. We note that these values are a few times higher than the previous $z\sim 0$ estimate of $n_{\rm SMBH}$ estimated by Vika et al. (2009) and Banik et al. (2019). At the highest redshifts, i.e., $z\sim 7-8$, we find $n_{\rm SMBH,raw-fp}\sim 10^{-4}\>{\rm cMpc}^{-3}$. In between these redshift extremes we derive smoothly declining intermediate values of $n_{\rm SMBH,raw-fp}$ as redshift increases.
@@
Our approach does not account for sources of incompleteness arising from the diverse AGN selection methods employed to compile the Lyu et al. (2022) sample. For instance, dust-obscured AGNs might remain undetected in X-ray surveys, or inadequate spectral resolution could hinder precise AGN diagnostics. Additionally, host galaxy contamination, which varies with redshift and host galaxy luminosity, could influence the number of variable sources recovered via variability studies, as well as the observed inverse relationship between variability significance and AGN luminosity (e.g., Suberlak et al., 2021). Nevertheless, we argue that incorporating these complexities would statistically increase the expected number of recovered AGNs in the field, of which only a fraction would be varying in the optical, consequently leading to lower recovered fractions (i.e., larger incompleteness corrections). In this scenario, our $F_{\rm var}$ would thus represent, despite considerable uncertainties, a conservative estimate. More extensive, longer term monitoring of AGN variability across a range of luminosities, wavelengths and redshifts is needed for improved estimates of $F_{\rm var}$.
@@
After the above correction for variability incompleteness, in the low-$z$ interval we obtain co-moving number densities of SMBHs of $n_{\rm SMBH,var}\simeq 2\times 10^{-2}\>{\rm cMpc}^{-3}$ (open diamonds in Fig. 6), with little sensitivity to the choice of significance level. This may indicate our estimate of the variability incompleteness factor as a function of significance level is reasonable. The smallest uncertainties result from the $2.5\sigma$ selected sample, formally yielding $n_{\rm SMBH,var}=1.7\pm 1.4\times 10^{-2}\>{\rm cMpc}^{-3}$. We regard this estimate as our best measure of the true local number density of SMBHs. However, it is likely to still be a lower limit if there are very faint AGN whose flux and flux variations would not be detected by the HST observations. At the highest redshifts, z=78z=7-8, we find $n_{\rm SMBH,var}\simeq 2\times 10^{-4}\>{\rm cMpc}^{-3}$. Given that we have adopted a redshift independent variability correction factor, the redshift dependence of $n_{\rm SMBH,var}$ is the same as that of $n_{\rm SMBH,raw-fp}$.
@@
Finally, we correct for “luminosity incompleteness” via a luminosity correction factor, $F_{\rm lum}$, i.e., $n_{\rm SMBH,lum}=F_{\rm lum}n_{\rm SMBH,var}$. In particular, at high redshift we are only able to detect relatively bright AGN (e.g., at $z\sim 6-7$ this corresponds at $M_{\rm UV}\simeq-18.6$), so this correction factor will boost the inferred number density of SMBHs. Following the method of Paper I and Harikane et al. (2023), who detected AGN via broad emission lines, we extrapolate from our detected AGN down an assumed luminosity function to a level of $M_{\rm UV}=-17\>$mag. We note that this is an arbitrary level, but by adopting this choice we are able to make a fair comparison to these previous studies. The value of the luminosity correction factor, $F_{\rm lum}$, generally becomes larger with redshift, but also depends on the form of the UV luminosity function (UVLF) we assume. Specifically, we integrate the double power-law fits detailed in Finkelstein & Bagley (2022) from redshift $z\sim 3$ to 9. At lower redshift, we integrate Schechter function fits to the UVLF from UVCANDELS and GALEX survey programs (Arnouts et al., 2005; Sun et al., 2024). At $z\lesssim 4$ we see that $F_{\rm lum}$ becomes smaller than unity as our detection limit becomes fainter than the UVLF lower bound of $M_{\rm UV}=-17$. Note that in the low $z$ interval, i.e., at z=00.5z=0-0.5, our sensitivity corresponds to an absolute magnitude of about $M_{\rm UV}=-8.5$ assuming an average redshift of $z\sim 0.25$. This explains why we consider $n_{\rm SMBH,var}$ as our best estimate at low redshifts. In Table 4 we report the values of $F_{\rm lum}$ used in each redshift interval.
@@
After the above correction factor for luminosity incompleteness, in the high-redshift regime (z=78z=7-8) we find $n_{\rm SMBH,lum}\sim 3\pm 6\times 10^{-3}\>{\rm cMpc}^{-3}$. We caution that this result is consistent with a null density and involves a relatively large luminosity function correction factor of $F_{\rm lum}\simeq 13$ and that the systematic uncertainty associated with this correction factor is not included in the above estimate. The redshift 6 to 7 interval has similarly high values of $n_{\rm SMBH,lum}$ (although the sample based on $2\sigma$ significance does not yield a number of sources greater than those expected from false positives). At intermediate redshifts, z=36z=3-6 we notice that $n_{\rm SMBH,lum}$ takes smaller values $\sim 10^{-3}-10^{-4}\>{\rm cMpc}^{-3}$. We do not expect such a rapid decline in the true number density of SMBHs, so this result may either indicate that a larger fraction of SMBHs were brighter than $M_{\rm UV}=-17\>$mag at $z>6$ or that the values at high redshift are somewhat overestimated via their luminosity function correction factor. At low redshifts, we see $n_{\rm SMBH,lum}$ rises again to values of a $\lesssim 10^{-3}\>{\rm cMpc}^{-3}$. However, at z=00.5z=0-0.5 this is a factor of about 20 smaller than the number we infer after variability completeness correction. This is simply a reflection of the fact that the observations are able to detect AGN that are much fainter than the arbitrary limit of $M_{\rm UV}=-17\>$mag.
@@
In Figure 7 we show our $3\sigma$ results for $n_{\rm SMBH,var}$ and $n_{\rm SMBH,lum}$, along with several other estimates of $n_{\rm SMBH}$. As mentioned, the previous estimate of $n_{\rm SMBH}(z=0)$ of $\sim 5\times 10^{-3}\>{\rm cMpc}^{-3}$ (Banik et al., 2019) and $\sim 9\times 10^{-3}\>{\rm cMpc}^{-3}$ (Vika et al., 2009) now appear to be superseded by about a factor of a few with our HUDF variability study estimate of $n_{\rm SMBH,var}$. Similarly, we see that our values of $n_{\rm SMBH,var}$ out to $z=3$ are about 2 times higher than the values of $n_{\rm SMBH}$ implied by the 31 previously known AGN in the HUDF (Lyu et al., 2022).
@@
At high redshifts we find consistency with our result from Paper I based on three detected AGN candidates between $z=6$ and 7. Our results for $n_{\rm SMBH,lum}$ at $z>6$ exceed those inferred from the JWST selected sample of Harikane et al. (2023) by about a factor of 5 From $z=4$ to 6 our estimates of $n_{\rm SMBH,lum}$ are instead about 5 times lower than those of Harikane et al. (2023). For all these high-$z$ results, there are large uncertainties due to the small number of directly detected sources. Indeed our $2\sigma$ selected sample at $z>7$ has the smallest Poisson errors.
@@
$\sigma 22$2 191 185 521 15/31 2.07
$2.5\sigma$ 92 67 68 188 11/31 2.81
$3\sigma$ 43 34 39 99+10 10/31 3.10
@@
Figure 7: As Fig. 6, but now only showing $n_{\rm SMBH}$ estimated from $3\sigma$ variables. Various other observational constraints are displayed: the red diamond shows the observational constraint derived in Paper I (Hayes et al., 2024) based on three detected AGN between $z=6$ and 7. Broad emission line sources at z=47z=4-7 (Harikane et al., 2023) are shown with the black diamonds, which use similar luminosity correction factors as our study. Brown stars indicate the number densities from the sample of 31 previously known AGN in the HUDF (Lyu et al., 2022). The black square shows a previous estimate of the local ($z=0$) co-moving number density of SMBHs (Banik et al., 2019).
@@ 3.4 Implications for SMBH Seeding Models
Our estimates of $n_{\rm SMBH}$ shown in Figures 6 and 7 place strong constraints on SMBH seeding models. At high redshifts, $z>6$, we see that our estimates of $n_{\rm SMBH,lum}\gtrsim 10^{-3}\>{\rm cMpc}^{-2}$ are about a factor of 10 higher than the fiducial Halo Mass Threshold (HMT) model used in the Illustris family of simulations (Vogelsberger et al., 2014) in which dark matter halos with masses $7\times 10^{10}\>M_{\odot}$ are seeded with SMBHs. The range of the green band shown in Figs. 6 and 7 corresponds to varying this threshold mass by factors of two. Thus an HMT model with a threshold mass of $\sim 3\times 10^{10}\>M_{\odot}$, which would be consistent with the $z=0$ constraints, would be closer to the high-$z$ data, but still low by about a factor of few.
Figures 6 and 7 also compares our constraints with the Romulus25 simulation (Tremmel et al., 2017). We see that this simulation, which seeds SMBHs using simple threshold conditions of gas properties (see §1), produces a high number density of SMBHs at early times, i.e., $n_{\rm SMBH}\sim 5\times 10^{-2}\>{\rm cMpc}^{-3}$ at $z\sim 12$. This rises by almost an order of magnitude by $z\simeq 4$, before then undergoing a modest decline by $z=0$. Given its high number densities of SMBHs, this model remains consistent with our observational constraints on $n_{\rm SMBH}$.
As discussed in §1, Direct Collapse (DC) seeing models struggle to reproduce the overall number densities of SMBHs. The most optimistic of the DC models discussed in §1 is that of Chon et al. (2016) with $n_{\rm SMBH}\sim 10^{-4}\>{\rm Mpc}^{-3}$. As shown in Figs. 6 and 7 this is about a factor of 50 below $n_{\rm SMBH}$ inferred from our HUDF variability study. Furthermore, the later simulations of DC by Wise et al. (2019) and Latif et al. (2022) find values of $n_{\rm SMBH}\lesssim 10^{-6}\>{\rm Mpc}^{-3}$, i.e., at least a factor of at least $10^{3}$ below our high-$z$ estimate. The constraints are even more severe when considering our low-$z$ estimates of $n_{\rm SMBH}$, which have smaller uncertainties and values of $\simeq 2\times 10^{-2}\>{\rm Mpc}^{-3}$.
@@
where $t_{*,10}=t_{*}/10\>$Myr, $t_{*}$ is the lifetime of the supermassive star, $_{53}\equiv S/10^{53}\>{\rm s}^{-1}, SS$ is the H-ionizing luminosity of the star, and $R_{R}$ is the radius of R-type HII region expansion achieved in the intergalactic medium. This estimate can be viewed as an upper limit since it has assumed maximal close packing of the sources and ignored other contributions to the ionizing background from Pop III.2, Pop II and AGN sources. It is interesting that our estimate of $n_{\rm SMBH}$ at low-$z$ from the HUDF is within a factor of a few of this fiducial estimate.
@@ 3.5 Comparison with other AGN diagnostics
Despite recent advances, it remains immensely challenging to obtain an accurate census of AGN at $z\gtrsim 6$ at faint magnitudes. Hot dust tracers (e.g., Stern et al., 2005) are shifted to the FIR, and X-ray and radio facilities are not sufficiently sensitive to identify individual faint AGN at these distances. High ionization UV emission lines (e.g., He ii $\lambda$1640, C iv $\lambda$1550, N v $\lambda$1240) are sometimes observed in luminous targets (e.g., Mainali et al., 2017; Laporte et al., 2017), but they can be inconclusive, as these lines also form in low-metallicity starbursts (Senchyna et al., 2020; Berg et al., 2019; Saxena et al., 2020). JWST-NIRSpec micro-shutter array (MSA) observations have delivered deep IR spectroscopy and several studies (e.g., Cameron et al., 2023; Boyett et al., 2024; Hu et al., 2024) could identify AGN via line flux ratios classification (BPT-like diagrams, Baldwin et al., 1981), although the targeted approach of MSA spectroscopy is subject to photometric pre-selection biases and tends to be focused more towards the brighter end of the luminosity function. In addition, NIRSpec surveys face challenges in achieving completeness at $M_{UV}\simeq-17$ to 20-20, where galaxies are $\sim 100$ times more abundant.